OpenStudy (anonymous):

suppose f(-2)=0, F(0)=pi, f(2)=2pi, and f'(x)=sqroot(4-x^2) if h(x)=f(2sinx) what is h'(0) i got 2cospi what i did was i took the derivative of 2sinx which is 2cosx and then it says the output value for x is pi so my answer is 2cospi which i think is wrong

OpenStudy (anonymous):

why would it be product rule

OpenStudy (anonymous):

i thought 2sinx was 1 term

OpenStudy (anonymous):

where did the f(x) come from?

OpenStudy (anonymous):

oh i think i know what you are talking, it wouldnt it be the chain rule then

OpenStudy (anonymous):

i viewed it as function f of x at 2sinx

OpenStudy (anonymous):

im given what i wrote so your guess is as good as mine :D

OpenStudy (anonymous):

crap I mean h'(0) = pi

OpenStudy (anonymous):

i know i hate it

OpenStudy (anonymous):

i will most deffinatly ask that question but you have helped me a lot

OpenStudy (anonymous):

h(x)=f(2sinx) h(x) = f(x) when x is equal to 2sin(x)

OpenStudy (anonymous):

its not the composition of two functions sorry

OpenStudy (anonymous):

My brain has really been failing me lately ugh Here is my try at this problem but I'm not sure if I'm right h'(x) =f'(2sin(x)) so h'(x) = (4-(2sin(x)')^2)^(1/2)' h'(x) = (4-(2cos(x))^(2))^(1/2)' h'(x) = (4+4cos(x)sin(x))^(1/2)' h'(x) = 1/2(4+4cos(x)sin(x)) * (-4sin(x)sin(x) + 4cos(x)cos(x)) h'(0) = (1/2(4)) * (4) h'(0) = 4/8 = 2

OpenStudy (anonymous):

where did that sinx come from?

OpenStudy (anonymous):

chain rule?

OpenStudy (anonymous):

Here is my logic if h(x) = f(2sinx) f(2sin(x)) = h(x) than f(2sin(x)) rearranges in a way that makes it equal to h(x) so if you take the derivative of 2sin(x) and place it into the know f'(x) function you should end up with a derivative equal to h(x)

OpenStudy (anonymous):

I don't know if that makes sense might be failed logic probably best you ask your ta or check your notes

OpenStudy (anonymous):

its sound logic to me

OpenStudy (anonymous):

I wouldn't take my word for it but that is just my reasoning and how I would answer it on a test if I had enough time lol after reading it incorrectly

OpenStudy (anonymous):

when you took the derivative did you subtract 1 from the power?

OpenStudy (anonymous):

I think I'm wrong actually I'm pretty sure im wrong

OpenStudy (anonymous):

Just ask your profs :L